### v3

@@ Abstract
With the growing application of computer vision in agriculture, image analysis has become essential for tasks such as crop health monitoring and pest detection. However, significant domain shifts caused by environmental variations, different crop types, and diverse data acquisition methods hinder model generalization across regions, seasons, and complex agricultural settings. This paper investigates how Domain Adaptation (DA) techniques can address these challenges by improving cross-domain transferability in agricultural image analysis. Given the limited availability of labeled data, weak model adaptability, and dynamic field conditions, DA has emerged as a promising solution. The review systematically summarizes recent advances in DA for agricultural imagery, focusing on applications such as crop health monitoring, pest detection, and fruit recognition, where DA methods have enhanced performance across diverse domains. DA approaches are categorized into shallow and deep learning methods, including supervised, semi-supervised, and unsupervised strategies, with particular attention to adversarial learning-based techniques that have demonstrated strong potential in complex scenarios. In addition, the paper reviews key public agricultural image datasets, evaluating their strengths and limitations in DA research. Overall, this work offers a comprehensive framework and critical insights to guide future research and development of domain adaptation in agricultural vision tasks.
@@ 1 Introduction
In agricultural image analysis, the high heterogeneity of data acquisition environments and the extreme scarcity of annotated data severely limit the generalization capability of models across different application scenarios. Among these challenges, the domain shift problem—referring to distribution discrepancies between source and target data—has emerged as a core obstacle hindering the deployment of agricultural intelligent vision systems [1, 2]. To address this, domain adaptation (DA) techniques have been widely introduced into agricultural image analysis in recent years, aiming to enhance model robustness and transferability across cross-regional, cross-temporal, and cross-device data.
@@
Agricultural image data acquisition typically relies on multiple sensors (e.g., RGB cameras, near-infrared cameras, hyperspectral sensors) and is susceptible to complex environmental factors such as illumination variations, seasonal changes, crop varieties, and soil backgrounds. These differences lead to significant distribution shifts for the same agricultural task across datasets, causing models trained under conventional paradigms to suffer drastic performance degradation in new environments [3]. For instance, models trained in laboratories or demonstration farms often exhibit sharp accuracy drops when deployed in actual fields, different regions, or across growing seasons.
@@
Moreover, annotating agricultural images requires expertise from agronomists and plant protection specialists, resulting in high costs and limited scalability. Compared to general computer vision tasks, the availability of high-quality annotated datasets in agriculture remains extremely scarce [4]. Consequently, directly transferring pre-trained models from natural image domains proves ineffective, underscoring the necessity of domain adaptation to reduce annotation dependency and enhance cross-domain adaptability. The core of domain adaptation lies in minimizing feature distribution discrepancies between source and target domains, enabling models to achieve comparable performance on the target domain.
Recently, DA techniques have demonstrated significant efficacy in key tasks such as crop disease/pest detection, hyperspectral image processing, crop growth monitoring, and precision farming management. Particularly under unsupervised or weakly supervised conditions, their ability to model data distribution shifts positions them as a critical pathway toward scalable agricultural visual intelligence. Although numerous surveys on transfer learning and domain adaptation exist in natural image analysis [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], systematic reviews tailored to agricultural image analysis remain insufficient. Building upon this, this survey aims to systematically review and summarize the current state of domain adaptation methods in agricultural image analysis. We categorize existing methods into two major classes (shallow and deep learning models) and discuss their supervised, semi-supervised, and unsupervised strategies. Particular focus is placed on adversarial learning-based adaptation methods and their responses to challenges such as seasonal variation, multi-source fusion, and annotation scarcity, ultimately offering a methodological reference for future intelligent agricultural systems.
@@ 2.1 Domain Shift in Agricultural Image Analysis
In agricultural image analysis, domain shift refers to the discrepancy in data distribution between the training dataset (source domain) and the test dataset (target domain). This shift can lead models to perform well on the source domain but fail to generalize effectively to the target domain, thereby significantly degrading overall model performance. Agricultural images often exhibit substantial distributional differences due to a variety of factors, including environmental variations, crop types, and imaging conditions. Specifically, images captured in different geographic regions or under varying environmental settings—such as changes in lighting, climate, seasonal dynamics, crop varieties, and camera viewpoints—exacerbate domain shift and present serious challenges to model transferability.
As shown in Figure 1, numerous studies have explored the use of domain adaptation (DA) techniques to address these challenges in agricultural vision tasks. DA methods aim to mitigate domain shift by aligning feature distributions between source and target domains, enabling more robust and generalizable model performance.
Figure 1: Illustration of source and target data: (left) the original feature distribution shows misalignment due to domain shift; (right) improved feature alignment after applying domain adaptation methods.
For example, in crop classification tasks, images collected from different regions often exhibit marked heterogeneity in appearance. Figure 2 illustrates this phenomenon using aerial scenes from Ghaziabad, India, captured during the rainy and dry seasons. The images clearly reflect substantial environmental differences, such as lighting and climate, that introduce distributional discrepancies between source and target domains. These shifts significantly impact the accuracy and robustness of key agricultural vision tasks, including crop monitoring and pest detection, particularly in cross-regional, cross-seasonal, and cross-device scenarios.
As the demand for scalable agricultural image analysis continues to grow, domain adaptation has emerged as a critical technology for improving cross-domain generalization. By reducing distributional gaps between domains, DA techniques enhance the reliability of agricultural models across diverse settings, ultimately supporting more accurate, consistent, and transferable performance in real-world agricultural applications.
@@ 2.2 Domain Adaptation and Transfer Learning
In a typical transfer learning (TL) framework, two fundamental concepts are involved: domain and task. A domain refers to a specific feature space along with its marginal probability distribution, while a task is defined by a label space and an associated predictive function. The goal of transfer learning is to leverage knowledge learned from a source task $T_{A}$ in domain $A$ and apply it to a target task $T_{B}$ in domain $B$. Notably, transfer learning allows for either the domain or the task—or both—to differ between the source and target settings.
Domain Adaptation (DA) is a specialized and widely studied subfield of transfer learning. In DA, the feature space and task remain consistent across domains, but the marginal distributions differ, i.e., $P_{S}(X)\neq P_{T}(X)$. Formally, let $X\times Y$ denote the joint feature and label space. The source domain $S$ and target domain $T$ are defined over this space, but their respective distributions $P_{S}$ and $P_{T}$ differ. Assume there are $n_{S}$ labeled samples in the source domain, $D_{S}=\{(x_{i}^{S},y_{i}^{S})\}_{i=1}^{n_{S}}$, and $n_{T}$ samples (either labeled or unlabeled) in the target domain, $D_{T}=\{x_{j}^{T}\}_{j=1}^{n_{T}}$.
The objective of domain adaptation is to transfer knowledge from the source domain $S$ to improve model performance on the target domain $T$ for a shared task. This is typically achieved by aligning the distributions of the two domains—either explicitly or implici
@@ 2.3 Domain Adaptation Problem Setup
In the agricultural domain, domain adaptation (DA) methods face unique challenges stemming from pronounced discrepancies between source and target domains. These discrepancies are particularly evident in three dimensions: spatial distribution, temporal sequences, and modality diversity. Agricultural image data are typically collected from heterogeneous platforms, including satellite remote sensing, unmanned aerial vehicles (UAVs), and ground-based sensors. Due to variations in climate conditions, crop types, and sensor characteristics across different environments, substantial distributional heterogeneity arises among agricultural datasets. To address this, domain adaptation techniques have become essential tools for enhancing model generalization by bridging distribution gaps across domains. Depending on the availability of labeled data in the target domain, DA methods are commonly categorized into supervised, semi-supervised, and unsupervised approaches, with unsupervised domain adaptation (UDA) drawing particular attention due to the scarcity and high cost of annotations in agricultural settings.
@@
Early efforts in domain adaptation primarily focused on aligning the data distributions between source and target domains through three main technical pathways: instance-based alignment (e.g., sample reweighting), feature space transformation (e.g., subspace learning), and classifier parameter adaptation. These traditional methods typically rely on hand-crafted features and statistical alignment techniques to reduce domain discrepancy. For instance, instance-based approaches adjust the contribution of each source sample to better match the target distribution; feature-based methods seek a shared subspace through linear projections; and classifier-based methods improve cross-domain performance by regularizing classifier parameters. These techniques, collectively referred to as shallow domain adaptation (Shallow DA), offer advantages such as low computational complexity, high training efficiency, and strong interpretability. Consequently, they are particularly well-suited for deployment in resource-constrained agricultural scenarios.
@@
The emergence of deep learning significantly reshaped the domain adaptation landscape. In 2014, Girshick et al. [17] introduced the Region-based Convolutional Neural Network (R-CNN), pioneering the use of convolutional neural networks (CNNs) for object detection. Building on this, Fast R-CNN [18] and later Faster R-CNN [19] integrated region proposal, feature extraction, and classification into a unified, end-to-end architecture, vastly improving both accuracy and efficiency. In the agricultural context, Fuentes et al. [20] applied the Faster R-CNN framework with VGGNet/ResNet backbones to detect and localize tomato diseases and pests, achieving a mean Average Precision (mAP) of 85.98% across 10 classes. This success exemplifies the widespread adoption of deep learning in agricultural image analysis. Owing to their reliance on deep feature extraction and complex architectures, these approaches fall under the umbrella of deep domain adaptation (Deep DA).
@@
In summary, domain adaptation methods in agricultural image analysis can be systematically divided into two major paradigms: shallow and deep adaptation. Shallow DA includes classical techniques based on instance reweighting, feature alignment, and classifier adjustment, while Deep DA encompasses modern methods that exploit deep neural networks and can be further categorized into unsupervised, semi-supervised, and supervised variants depending on target label availability (see Table I). These two paradigms offer complementary strengths in terms of computational cost, scalability, and domain transferability, together forming a comprehensive framework for addressing cross-domain challenges in agricultural visual perception.
@@
Feature-based SA, TSSA, ECMDCM, CropSTGAN, MultiCropGAN, Spatial-Invariant Features, Invariant Features in RKHS, Sensor-driven Hierarchical DA, PSO-TrAdaBoost, ESMSL, HSL-GM
@@
Deep DA Semi-supervised learning MAML, WheatSeedBelt, SSDA-WheatHead, SSDA-WheatSeg, TDA-YOLO, SCDAL, CDADA
Unsupervised learning MSFF, GAN-DA, Standard GAN, DeepDA-Net, U-DA-Net, TriADA, MaxEnt-DA, DAE-DANN, Self-Attn-DA, ADANN, OpenDA, AdaptSegNet, ADVENT, BDL, DFENet, TSAN, DAN, WDGRL, CORAL, Transformer-UDA, MRAN, WPS-DSA, NCADA, TCANet, DDA-Net, TDDA, TST net, AMF-FSL, MSUN, MSCN, AMRAN, DJDANs, CLA
@@ 3.1 Instance Weighting
Molina-Cabanillas et al. [22] introduced DIAFAN-TL (Domain-Invariant Agricultural Features Adaptation Network for Transfer Learning), an instance-weighted learning algorithm for phenological prediction in olive tree cultivars. This method dynamically assigns weights to training samples based on inter-domain relevance, thus enhancing information diversity while reducing sample inconsistency.
Yaras et al. [23] proposed Random Histogram Matching (RHM), which leverages data augmentation to improve satellite image domain adaptation. By modeling sensor differences and environmental variations as nonlinear pixel transformations, and applying deep neural networks to augment data, this method boosts robustness to domain shifts.
For further framework optimization, Cui et al. [24] developed the Iterative Weighted Active Transfer Learning (IWATL) approach for hyperspectral image (HSI) classification. Through a dual-evaluation mechanism, it adaptively updates source sample weights, significantly enhancing classification performance. Li et al. [25] proposed the Cost-Sensitive Self-Paced Learning (CSSPL) framework, which integrates hybrid weight regularization for selecting multi-temporal samples, addressing seasonal and temporal variances in agricultural monitoring.
@@ 3.2 Feature Transformation
Feature-based domain adaptation methods aim to map data from the source and target domains into a shared, domain-invariant feature representation space. This is typically achieved through techniques such as subspace alignment, manifold learning, and low-rank representation. The primary objective of these methods is to reduce the distribution discrepancy between domains during the process of joint feature extraction [26].
As illustrated in Figure 4, these methods first construct a common feature space that captures domain-invariant properties. Subsequently, a classifier is trained using the transformed source domain features and corresponding labels, enabling the prediction of target domain samples with improved accuracy. Implementation strategies generally fall into two categories: subspace-based projection reconstruction and nonlinear transformation-based feature re-encoding.
The overarching goal is to align the statistical properties of features across domains by transforming the feature space, thereby enhancing the generalization ability of the model on the target domain. This approach is particularly effective in agricultural scenarios characterized by domain shifts induced by seasonal, geographical, or sensor-based variations.

### v4

@@ Abstract
With the growing application of computer vision in agriculture, image analysis has become essential for tasks such as crop health monitoring and pest detection. However, significant domain shifts caused by environmental variations, different crop types, and diverse data acquisition methods hinder model generalization across regions, seasons, and complex agricultural settings. This paper investigates how Domain Adaptation (DA) techniques can address these challenges by improving cross-domain transferability in agricultural image analysis. Given the limited availability of labeled data, weak model adaptability, and dynamic field conditions, DA has emerged as a promising solution. The review systematically summarizes recent advances in DA for agricultural imagery, focusing on applications such as crop health monitoring, pest detection, and fruit recognition, where DA methods have improved performance in diverse domains. DA approaches are categorized into shallow and deep learning methods, including supervised, semi-supervised, and unsupervised strategies, with particular attention to adversarial learning-based techniques that have demonstrated strong potential in complex scenarios. In addition,the paper reviews key public agricultural image datasets, evaluating their strengths and limitations in DA research. In general, this work offers a comprehensive framework and critical insights to guide future research and development of domain adaptation in agricultural vision tasks.
@@ 1 Introduction
In agricultural image analysis, the high heterogeneity of data acquisition environments and the extreme scarcity of annotated data severely limit the generalizability of models across different application scenarios. Among these challenges, the domain shift problem—referring to distribution discrepancies between source and target data—has emerged as a key obstacle hindering the deployment of intelligent agricultural vision systems [1, 2]. To address this issue, domain adaptation (DA) techniques have recently gained prominence in agricultural image analysis, aiming to enhance model robustness and transferability across cross-regional, cross-temporal, and cross-device data sources.
@@
Agricultural image data are typically acquired using various sensors (e.g., RGB cameras, near-infrared imagers, hyperspectral sensors) and are highly susceptible to environmental factors such as illumination changes, seasonal cycles, crop varieties, and soil backgrounds. These differences lead to significant distributional shifts for the same task across datasets, causing models trained under conventional paradigms to suffer severe performance degradation when deployed in new environments [3]. For example, models trained on laboratory or demonstration farm data often exhibit sharp accuracy drops when applied to real-world fields, other geographic regions, or different crop growth periods.Domain adaptation extends beyond crops to livestock management, as seen in Dubourvieux et al.’s work on cattle re-identification [4]. Their cumulative multi-domain strategy addresses challenges like varying barn conditions or camera viewpoints, demonstrating DA’s versatility in animal farming.
@@
Furthermore, annotating agricultural images generally requires expert knowledge from agronomists and plant protection specialists, resulting in high annotation costs and limited data scalability. Compared to general computer vision tasks, the availability of high-quality labeled datasets in agriculture remains extremely limited [5]. Consequently, directly transferring pre-trained models from natural image domains proves ineffective, highlighting the necessity of DA techniques to reduce annotation dependency and improve cross-domain adaptability. The core objective of DA is to minimize feature distribution discrepancies between the source and target domains, enabling models to perform comparably in the target domain.
In recent years, DA has shown significant promise in key agricultural applications such as crop disease and pest detection, hyperspectral image analysis, crop growth monitoring, and precision farming management. Particularly under unsupervised or weakly supervised settings, the ability of DA to model distributional shifts makes it a vital enabler of scalable agricultural visual intelligence. While numerous reviews on transfer learning and domain adaptation exist in the field of natural image analysis [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], systematic reviews specifically focused on agricultural image analysis remain limited.
To address this gap, this review systematically summarizes the current landscape of domain adaptation methods in agricultural image analysis. We categorize existing approaches into two major classes—shallow and deep learning models—and examine their respective supervised, semi-supervised, and unsupervised strategies. Particular emphasis is placed on adversarial learning-based adaptation methods and their effectiveness in addressing key challenges such as seasonal variation, multi-source data fusion, and annotation scarcity. This review ultimately serves as a methodological reference for the development of next-generation intelligent agricultural systems.
@@ 2.1 Domain Shift in Agricultural Image Analysis
In agricultural image analysis, domain shift refers to the discrepancy in data distribution between the training dataset (source domain) and the test dataset (target domain). This shift can cause models to perform well on the source domain but fail to generalize effectively to the target domain, thereby significantly degrading overall performance. Agricultural images often exhibit substantial distributional differences due to various factors, including environmental conditions, crop types, and imaging configurations. Specifically, images captured in different geographic regions or under varying environmental settings—such as changes in lighting, climate, seasonal dynamics, crop varieties, and sensor viewpoints—can exacerbate domain shift, presenting serious challenges to model transferability.
As shown in Figure 1, numerous studies have explored the use of domain adaptation (DA) techniques to address these challenges in agricultural vision tasks. DA methods aim to mitigate domain shift by aligning feature distributions between the source and target domains, enabling more robust and generalizable performance.
Figure 1: Illustration of source and target data: (left) misaligned feature distribution due to domain shift; (right) improved alignment after applying domain adaptation.
For example, in crop classification tasks, images collected from different regions often exhibit marked heterogeneity in appearance. Figure 2 illustrates this phenomenon using aerial scenes from Ghaziabad, India, captured during the rainy and dry seasons. The images clearly reflect environmental differences, such as lighting and climate, which introduce significant distributional discrepancies. These shifts can impact the accuracy and robustness of key agricultural vision tasks, including crop monitoring and pest detection—especially in cross-regional, cross-seasonal, and cross-device scenarios.
As the demand for scalable agricultural image analysis continues to grow, domain adaptation has emerged as a critical technology for improving cross-domain generalization. By reducing distributional gaps between domains, DA techniques enhance the reliability of agricultural models across diverse settings, ultimately supporting more accurate, consistent, and transferable performance in real-world applications.
@@ 2.2 Domain Adaptation and Transfer Learning
In a typical transfer learning (TL) framework, two fundamental concepts are involved: domain and task. A domain refers to a specific feature space along with its marginal probability distribution, while a task is defined by a label space and an associated predictive function. The goal of TL is to leverage knowledge learned from a source task $T_{A}$ in domain $A$ and apply it to a target task $T_{B}$ in domain $B$. Notably, either the domain, the task, or both may differ between the source and target settings.
Domain Adaptation (DA) is a specialized and widely studied subfield of TL. In DA, the feature space and task remain consistent across domains, but the marginal distributions differ, i.e., $P_{S}(X)\neq P_{T}(X)$. Formally, let $X\times Y$ denote the joint feature and label space. The source domain $S$ and target domain $T$ are defined over this space, but their respective distributions $P_{S}$ and $P_{T}$ differ. Assume there are $n_{S}$ labeled samples in the source domain, $D_{S}=\{(x_{i}^{S},y_{i}^{S})\}_{i=1}^{n_{S}}$, and $n_{T}$ samples (either labeled or unlabeled) in the target domain, $D_{T}=\{x_{j}^{T}\}_{j=1}^{n_{T}}$.
The objective of domain adaptation is to transfer knowledge from the source domain $S$ to improve model performance on the target domain $T$ for a shared task. This is typically achieved by aligning the distributions of the two domains—either explicitly or implicitly—through various adaptation strategies.
@@ 2.3 Domain Adaptation Problem Setup
In the agricultural domain, domain adaptation methods face unique challenges stemming from pronounced discrepancies between source and target domains. These discrepancies are particularly evident in three dimensions: spatial distribution, temporal sequences, and modality diversity. Agricultural image data are typically collected from heterogeneous platforms, including satellite remote sensing, unmanned aerial vehicles (UAVs), and ground-based sensors. Due to variations in climate conditions, crop types, and sensor characteristics across different environments, substantial distributional heterogeneity arises. To address this, DA techniques have become essential for enhancing generalization by bridging distribution gaps across domains.
Depending on the availability of labeled data in the target domain, DA methods are commonly categorized into supervised, semi-supervised, and unsupervised approaches. Among them, unsupervised domain adaptation (UDA) is especially critical in agriculture due to the scarcity and cost of annotations.
@@
Early efforts in DA primarily focused on aligning data distributions between source and target domains through three main pathways: instance-based alignment (e.g., sample reweighting), feature space transformation (e.g., subspace learning), and classifier parameter adaptation. These traditional approaches typically rely on hand-crafted features and statistical alignment techniques to reduce domain discrepancy. For instance, instance-based approaches adjust the contribution of each source sample to better match the target distribution; feature-based methods seek a shared subspace through linear projections; and classifier-based methods improve cross-domain performance by regularizing classifier parameters. These methods, collectively referred to as shallow domain adaptation (Shallow DA), offer advantages such as low computational complexity, high efficiency, and strong interpretability—making them suitable for resource-constrained agricultural applications.
@@
The emergence of deep learning has significantly reshaped the DA landscape. In 2014, Girshick et al. [18] introduced the Region-based Convolutional Neural Network (R-CNN), pioneering CNNs for object detection. Fast R-CNN [19] and Faster R-CNN [20] further improved this framework by integrating region proposal, feature extraction, and classification into an end-to-end pipeline. In agriculture, Fuentes et al. [21] applied Faster R-CNN with VGGNet/ResNet backbones to detect and localize tomato diseases and pests, achieving a mean Average Precision (mAP) of 85.98% across 10 classes. This success exemplifies the widespread adoption of deep domain adaptation (Deep DA) in agricultural image analysis.
@@
In summary, DA methods in agriculture can be categorized into two paradigms: shallow and deep adaptation. Shallow DA includes classical techniques based on instance reweighting, feature alignment, and classifier adjustment, while Deep DA encompasses modern methods leveraging deep neural networks. These can be further classified into supervised, semi-supervised, and unsupervised approaches depending on label availability, as summarized in Table 1. Each paradigm offers complementary strengths in terms of computational efficiency, scalability, and generalization.
@@
Feature-based SA, TSSA, ECMDCM, CropSTGAN, MultiCropGAN, Spatial-Invariant Features, RKHS Features, Sensor-driven Hierarchical DA, PSO-TrAdaBoost, ESMSL, HSL-GM
@@
Deep DA Semi-supervised MAML, WheatSeedBelt, SSDA-WheatHead, SSDA-WheatSeg, TDA-YOLO, SCDAL, CDADA
Unsupervised MSFF, GAN-DA, DeepDA-Net, U-DA-Net, TriADA, MaxEnt-DA, DAE-DANN, Self-Attn-DA, ADANN, OpenDA, AdaptSegNet, ADVENT, BDL, DFENet, TSAN, DAN, WDGRL, CORAL, Transformer-UDA, MRAN, WPS-DSA, NCADA, TCANet, DDA-Net, TDDA, TST-Net, AMF-FSL, MSUN, MSCN, AMRAN, DJDANs, CLA
@@ 3.1 Instance Weighting
Molina-Cabanillas et al. [23] introduced DIAFAN-TL (Domain-Invariant Agricultural Features Adaptation Network for Transfer Learning), an instance-weighted learning algorithm for phenological prediction in olive tree cultivars. This method dynamically assigns weights to training samples based on inter-domain relevance, thereby enhancing information diversity while reducing sample inconsistency. Yaras et al. [24] proposed Random Histogram Matching (RHM), which leverages data augmentation to improve satellite image domain adaptation. By modeling sensor differences and environmental variations as nonlinear pixel transformations and applying deep neural networks for augmentation, this method improves robustness to domain shifts.
To further optimize the framework, Cui et al. [25] developed the Iterative Weighted Active Transfer Learning (IWATL) approach for hyperspectral image (HSI) classification. Through a dual-evaluation mechanism, it adaptively updates source sample weights, significantly enhancing classification performance. Li et al. [26] proposed the Cost-Sensitive Self-Paced Learning (CSSPL) framework, which integrates hybrid weight regularization for selecting multi-temporal samples, addressing seasonal and temporal variances in agricultural monitoring.
@@ 3.2 Feature Transformation
Feature-based domain adaptation methods aim to map data from the source and target domains into a shared, domain-invariant feature representation space. This is typically achieved through techniques such as subspace alignment, manifold learning, and low-rank representation. The primary objective is to reduce the distributional discrepancy between domains during the process of joint feature extraction [27].
As illustrated in Figure 4, these methods first construct a common feature space that captures domain-invariant properties. A classifier is then trained using the transformed source domain features and corresponding labels, enabling more accurate predictions for target domain samples. Implementation strategies generally fall into two categories: subspace-based projection reconstruction and nonlinear transformation-based feature re-encoding.
The overarching goal is to align the statistical properties of features across domains by transforming the feature space, thereby enhancing the model’s generalization ability in the target domain. This approach is particularly effective in agricultural scenarios characterized by domain shifts induced by seasonal, geographical, or sensor-based variations.